OpenStudy (anonymous):

18 percent of what number is 108

OpenStudy (blurbendy):

.18x = 108 solve for x

OpenStudy (bakonloverk):

pu IS over OF

OpenStudy (bakonloverk):

put*

OpenStudy (bakonloverk):

\[\frac{ 18%}{ 100} =\frac{ 108 }{ x }\] solve the proportion

OpenStudy (anonymous):

(18/100)x=108 x=108*100/18=3*3*2*6*100/3*3*2=600